Documenta Mathematica

Band 11, 2006

Igor Wigman Statistics of Lattice Points

in Thin Annuli for Generic Lattices  1–23

Haruzo Hida

Automorphism Groups of Shimura Varieties  25–56

Sandra Rozensztajn Compactification de Schémas Abéliens

Dégénérant au-dessus d’un Diviseur Régulier  57–71

W. Bley On the Equivariant Tamagawa Number Conjecture for Abelian Extensions

of a Quadratic Imaginary Field  73–118

Ernst-Ulrich Gekeler The Distribution of Group Structures

on Elliptic Curves over Finite Prime Fields  119–142

Claude Cibils and Andrea Solotar Galois coverings, Morita Equivalence

and Smash Extensions of Categories over a Field  143–159

Bernd Ammann, Alexandru D. Ionescu, Victor Nistor Sobolev Spaces on Lie Manifolds and

Regularity for Polyhedral Domains  161–206

Srikanth Iyengar, Henning Krause Acyclicity Versus Total Acyclicity

for Complexes over Noetherian Rings  207–240

Anton Deitmar and Joachim Hilgert Erratum to: Cohomology of Arithmetic Groups with Infinite Dimensional Coefficient Spaces

cf. Documenta Math. 10, (2005) 199–216  241–241

Matthias Franz

Koszul Duality and Equivariant Cohomology  243–259

Christian Böhning Derived Categories of Coherent Sheaves

on Rational Homogeneous Manifolds  261–331

Goro Shimura Integer-Valued Quadratic Forms

and Quadratic Diophantine Equations  333–367

Imre Bárány and Günter Rote

Strictly Convex Drawings of Planar Graphs  369–391

Achill Schürmann

On Packing Spheres into Containers  393–406

Philip Foth, Michael Otto A Symplectic Approach

to Van Den Ban’s Convexity Theorem  407–424

Henri Gillet and Daniel R. Grayson

Volumes of Symmetric Spaces via Lattice Points  425–447

Chia-Fu Yu The Supersingular Loci and Mass

Formulas on Siegel Modular Varieties  449–468

Dabbek Khalifa et Elkhadhra Fredj

Capacité Associeé a un Courant Positif Fermé  469–486